# ORFS-agent: Tool-Using Agents for Chip Design Optimization

### v1

@@ IV Experiments and Results
In our experiments, we use six design instances comprising three circuits (IBEX, AES and JPEG) in two technology nodes (SKY130HD and ASAP7). All experiments are performed on a Google Cloud Platform (GCP) virtual machine with 112 virtual CPUs (C2D AMD Milan) and 220GB RAM. The flow is based on the ce8d36a commit of the OpenROAD-flow-scripts repository, along with associated versions of OpenROAD, KLayout, Yosys, and other dependencies. [^1] To confirm reproducibility, we transfer the DEF files generated during the place-and-route process to another server and verify that the results are consistent across environments.
Although the timeout to run depends on the circuit, each “iteration” is consistently defined as 25 parallel runs, with the exception of the JPEG circuits, which are run with $12$ parallel runs each. Each of these parallel runs uses four vCPUs and 4 GB of RAM. These parallel runs write to different intermediate folders for logs and results, running Yosys and other steps in parallel without reusing intermediate steps. At the end of each iteration, a clean operation moves the logs and folders to a separate directory to avoid contamination.
Our experiments cover two settings. The first, “4-variable” setting involves optimizing over four parameters: Core Utilization, TNS End Percent, LB Add On Place Density, and Clock Period. The first three are in the configuration files of the respective circuits; Clock Period is in the .sdc file. The second, “12-variable” setting has eight additional tunable parameters that guide the physical design flow more comprehensively. We distinguish three experimental cases:
@@
  • 12-variable with tool use, which incorporates all 12 variables along with the optimization tools.
Our experimental choices reflect the LLM requiring more complex tools to deal with higher-dimensional optimization problems. All of our experiments involve either $375$ (4-variable setting) or $600$ iterations (12-variable setting). The ORFS-agent parameters are configured as follows: (i) the model used is claude-3-5-sonnet-20241022; [^2] We do not observe any performance gains from the o1-class of models. (ii) temperature is set to $0.1$; (iii) nucleus and top-K sampling are disabled; and (iv) to always output $25\times 4/12$ parameters (the number of parameters the LLM is optimizing over) in the specified format, we utilize JSON Mode (Structured Outputs API). [^3] In terms of API costs, all of our experiments, including all failed runs, cost USD \$48. Claude Sonnet 3.5 has since been overtaken by models which are simultaneously better and cheaper by 5-10$\times$ per token. As noted above, approaches such as ChatEDA require fine-tuning, which costs significantly more (by at least 10-20$\times$) and requires inference costs for serving the custom model. Appendix B provides additional experimental details.
@@
TABLE I: OR-AutoTuner results. Top block: 375 iterations with 4 parameters. Bottom block: 1000 iterations with 12 parameters. Columns are grouped by optimization: WL-only (WL, ECP*), ECP-only (ECP, WL*), and co-optimization (WL**, ECP**). Bold font indicates superior results.
@@
TABLE II: ORFS-agent results for various tuning settings. Top: 375 iterations, 4 parameters (no tool); Middle: 375 iterations, 4 parameters (with tool); Bottom: 600 iterations, 12 parameters (with tool). Each row reports the figures of merit under WL-only, ECP-only, and Co-optimization objectives. Bold font indicates superior results.
@@
  • $\mathbf{ECP}^{*}$ (resp. $\mathbf{WL}^{*}$) denotes the value of $\mathbf{ECP}$ (resp. $\mathbf{WL}$) when $\mathbf{WL}$ (resp. $\mathbf{ECP}$) is optimized as the objective.
  • $\mathbf{ECP}^{**}$ and $\mathbf{WL}^{**}$ represent the values obtained when both metrics are jointly optimized in a multi-objective setting.
We use OR-AutoTuner as the baseline. We run OR-AutoTuner for $1000$ iterations with $12$ parameters to optimize, and for $375$ iterations with $4$ parameters to optimize. Results are shown in Table I.
In the following, Section IV-A presents results for single-objective optimization. Then, Section IV-B gives results for multi-objective optimization. We then demonstrate ORFS-agent’s capability to perform constrained optimization guided by natural language commands. Appendix C provides additional ablation studies for the ASAP7 IBEX design. Studies on data leakage, model confidence and robustness are presented in Appendix D.
@@ IV-A Single-objective Optimization
There is only a single objective to optimize in this problem. You observe that the value inside the final JSON is $\mathbf{WL}$, while the corresponding baseline value is $\mathbf{WL_{\alpha}}$. Your effective loss to minimize should be $\mathbf{\frac{WL}{WL_{\alpha}}}$.”
where $\mathbf{WL_{\alpha}}$ denotes the routed wirelength obtained using the default configuration of OpenROAD-flow-scripts and serves as a fixed baseline for normalization. [^6] The baseline results obtained using the default configuration of OpenROAD-flow-scripts are shown in Appendix IV-C.
However, due to runtime timeouts, typically occurring during the detailed routing stage, the routed wirelength $\mathbf{WL}$ may not always be available. In contrast, the estimated wirelength at the end of CTS $\mathbf{WL^{\prime}}$ almost always exists. If $\mathbf{WL}$ is absent but $\mathbf{WL^{\prime}}$ exists, we include the following instruction:
You do not observe the correct value $\mathbf{WL}$, but you observe a strong surrogate $\mathbf{WL^{\prime}}$, which you should use as a signal. Your effective loss should be $\mathbf{\frac{WL^{\prime}}{WL^{\prime}_{\alpha}}}$.
where $\mathbf{WL^{\prime}_{\alpha}}$ denotes post-CTS wirelength obtained under the same default configuration. This surrogate-based guidance enables the LLM to continue learning from partial information while maintaining alignment with the optimization objective.
The experimental results are shown in Table II. Before discussing these results, we would like to highlight that our results are obtained with the Claude 3.5 class of models.
@@
  • Comparison between default ORFS and ORFS-agent: ORFS-agent (600 iterations, 12 parameters with tool use) achieves an average improvement of $13.2\%$ in wirelength (or ECP) when configured for wirelength (or ECP) optimization.
  • Comparison between OR-AutoTuner and ORFS-agent: Despite using 40% fewer iterations, ORFS-agent (600 iterations, 12 parameters, with tool use) achieves better results than OR-AutoTuner on $12$ out of $24$ metrics measured for single-objective optimization, thereby exactly drawing on cases of superiority.
@@
  • We notice that ORFS-agent does not significantly worsen the metric not being optimized. For instance, when optimizing wirelength on ASAP7 IBEX, the ECP obtained by ORFS-agent (600 iterations, 12 parameters with tool use) is $40\%$ of that obtained by OR-AutoTuner. We hypothesize this is from the ECP values entering the agent’s context window, causing it to avoid actions that would worsen ECP.
Figure 3: Optimization trajectory of ASAP7 IBEX (co-optimization, 600 iterations, 12 parameters with tool use). Wirelength (WL), effective clock period (ECP), instance count (Count) are normalized to baseline values from default ORFS settings.
@@ IV-B Multi-objective Optimization
In the case of multiple losses—e.g., $\mathbf{WL}$ and $\mathbf{ECP}$—we instruct the LLM to consider the loss:
$\mathbf{\frac{WL}{WL_{\alpha}}}+\mathbf{\frac{ECP}{ECP_{\alpha}}}$ (1)
with the same substitution, if applicable, for $\mathbf{ECP^{\prime},WL^{\prime}}$ if any of these are unavailable, with the same phrasing adopted above. The experimental results are shown in Table II. Key results are as follows:
  • Sustained gains. Unlike single-objective optimization, where gains observed exceed $13\%$, gains on multi-objective cases are roughly $11.3\%$ on ECP and $10.5\%$ on wirelength simultaneously. This is measured relative to the baselines discussed in subsection IV-C, having $12$ parameters to optimize for $600$ iterations with tool use. Out of $12$ comparisons to OR-AutoTuner, ORFS-agent is superior in $6$ of them, thereby exactly drawing.
  • No worsening. As with single-objective optimization, ORFS-agent does not worsen any metric - either ECP or wirelength, relative to baselines. In fact, both QoRs improve simultaneously.
We visualize the optimization trajectory of instance count alongside wirelength and ECP in Figure 3. We notice that instance count can actually improve for free. To better understand gains on auxiliary metrics (metrics that do not directly enter the optimization function), we examine correlational properties of metrics in Figure 4. We look at how other metrics vary in an optimization run as we carry out ORFS-level runs for ASAP7 IBEX under different parameters and differing optimization goals - namely wirelength, ECP, or a combination of the two as defined above. Some metrics do not “come for free” - ECP and power, for instance, have an inverse relationship with each other.
Figure 4: Correlation of metrics for ASAP7 IBEX: wirelength (WL), effective clock period (ECP), instance area (Area), instance count (Instances) and power (Power). Pearson correlation captures linear relationships; Kendall correlation captures rank-based relationships.
@@ IV-C Baseline comparisons
We obtain baseline solution metrics using the default ORFS flow parameters seen at the cited commit hash. These metrics are shown in Tables III and  IV. Compared to these baselines, ORFS-agent achieves an improvement of over 13% versus the default ORFS flow parameters, when averaged across all six testcases and single-metric (i.e., ECP and wirelength) comparisons. (For all metrics, lower is better.) For additional context, we also provide in Table V the post-route metrics obtained by a proprietary, closed-source tool on ASAP7 circuits; these are quoted from [21]. [^7] Via correspondence with authors of [21], we understand that OpenROAD uses less memory and vCPU resource than the commercial tool, by factors of 5$\times$ or more.
CTSWL CTSECP DRWL ECP
@@
TABLE III: Parameter values for original circuits, no changes.
@@
TABLE IV: Derived values for original circuits.
@@ IV-D Constrained Optimization under Natural Language Commands
A human engineer is typically given commands significantly different from the utility functions used by OR-AutoTuner. Specifically, a common class of optimization commands relayed to humans takes the form: “Optimize metric $X$ as much as possible, while not allowing metrics $A,B,C$ to deteriorate by more than $2\%$.” This percentage value is with reference to the previously mentioned baseline values in Tables III and  IV. This form of constrained optimization is generally not straightforward with existing OR-AutoTuner setups. LLMs can attempt such optimization by incorporating these constraints directly into their context. We study this phenomenon in Table VI, where we optimize $ECP$ and $WL$ subject to a maximum worsening of $2\%$ and $4\%$ in other metrics of interest—Area, Instance Count, Power, and PDP. We term the maximum worsening allowed as “leeway”. We observe that this often improves multiple metrics simultaneously. This avoids the problem of, e.g., worsening power values that we observed in Fig 4. Comparisons to OR-AutoTuner and further discussion appears in the appendix, alongside tests for robustness and prompt-level sensitivity. ORFS-agent respects the constraints (in terms of maximum $2\%$ or $4\%$ worsening) imposed. Simultaneously, it improves the metrics to be optimized relative to baseline values provided previously.
@@
TABLE VI: Constrained optimization results. Red positive numbers indicate worsening, i.e., an increase relative to baseline, and the “leeway” indicates the maximum worsening allowed.
@@ V Conclusion and Future Directions
In this work, we propose ORFS-agent, a plug-in open-source framework that integrates LLMs into OpenROAD for hyperparameter optimization. Despite using 40% fewer iterations, ORFS-agent achieves results comparable to OR-AutoTuner. ORFS-agent adopts a modular, provider-agnostic framework, which makes it user-friendly and easy to adopt. ORFS-agent can switch seamlessly among LLM providers (e.g., Anthropic, OpenAI, Gemini) without reconfiguration. Additionally, ORFS-agent avoids the cost and complexity of fine-tuning and deployment, and requires no retraining or infrastructure changes when a new LLM is released. Thus, our ORFS-agent can instantly benefit from model breakthroughs. Unlike fine-tuning-based approaches, ORFS-agent can rapidly and cost-effectively adopt stronger models, making it both competitive and appealing.
Our future work will explore the following directions. (i) The field has seen the emergence of reasoning-based models, which use a longer chain of thought process before making decisions. These reasoning-based models are then fine-tuned to task-specific agentic models. We aim to integrate such agentic models into ORFS-agent to enhance its decision-making depth and adaptability. (ii) Our current framework uses a fixed set of pre-defined tools. In future work, we will enable the LLM agent to generate its own tools dynamically, which avoids the hand-coding of specific tools and increases the flexibility in handling different circuits and technology nodes. (iii) We plan to integrate ORFS-agent with search APIs and RAG tools, which enable the agent to access documentation, literature or previous published results. And (iv) we will evaluate ORFS-agent on larger-scale circuits and with more hyperparameters to assess its scalability and robustness. In combination with open-sourcing and OpenROAD integration, we believe that this work will add to the foundations for new research on LLMs for physical design and EDA.
@@ Appendix A EDA Flow and OpenROAD
The EDA flow is highly iterative , and suboptimal outcomes at one step may require going back to earlier steps for re-optimization (e.g., re-synthesizing netlist or adjusting placement). Each step generates logs and metrics that can be integrated into ML pipelines to guide subsequent decisions.
Alongside commercial tools, the OpenROAD project [3] provides a fully open-source, end-to-end RTL-to-GDS solution. It integrates logic synthesis, floorplanning, placement, clock tree synthesis and routing into a fully automated framework. OpenROAD is particularly well-suited for ML research due to the following advantages:
  • Transparency and Instrumentation: Researchers have direct access to internal data structures and logs, enabling detailed data collection and analysis.
  • Reproducibility: The open-source nature of OpenROAD facilitates replication by third parties.
  • Extensibility: APIs are available for integrating custom ML modules. For instance, OR-AutoTuner [22] applies Bayesian optimization (BO) to tune synthesis and place-and-route (SP&R) recipes.
OpenROAD provides a powerful sandbox for exploring data-driven approaches through the digital implementation process. Its open-source foundation enables researchers to modify the details of each implementation step and apply advanced ML techniques for parameter tuning, downstream prediction and flow optimization. In contrast, commercial tools are nearly unusable for open-source academic ML research, due to extremely high license fees and legal constraints on sharing even partial outputs.
Problems with Working with Proprietary Tools. Commercial EDA tools from industry leaders like Synopsys and Cadence are integral to modern chip design, offering comprehensive solutions that cover the entire design flow, including logic synthesis, simulation, and verification. These tools are renowned for their robustness and efficiency, making them the go-to choice for many semiconductor companies.
However, for machine learning (ML) practitioners and researchers, these proprietary tools present several challenges:
  • Closed Architectures: The proprietary nature of these tools means their internal workings are not accessible. This opacity hinders the ability to customize or extend functionalities, which is often necessary for integrating novel ML algorithms or for conducting in-depth research that requires toolchain modifications. Indeed, even the partial outputs cannot be shared without violating legal constraints, a strong contrast from the open-source ecosystem of Machine Learning.
  • High Costs: Licensing fees for commercial EDA tools can be prohibitively expensive, especially for academic institutions, startups, or individual researchers operating under budget constraints. This financial barrier limits the accessibility of these tools for experimental and exploratory ML applications.
  • Limited Flexibility: The closed file structures and lack of transparency in proprietary tools make it difficult to add customized features or promptly address bugs. This rigidity can impede the iterative and experimental nature of ML research, where rapid prototyping and testing of new ideas are crucial.
These constraints underscore the need for open-source alternatives that provide the flexibility, transparency, and cost-effectiveness required for ML-driven EDA research and development.
@@
Open-Source EDA and OpenROAD as an ML Testbed. To systematically study LLM- and ML-driven EDA optimizations, we focus on the OpenROAD project [3]. In contrast to commercial EDA offerings, OpenROAD is fully open-source, thus avoiding copyright restrictions and enabling access to the internal details of its place-and-route algorithms, log files, and data structures. This open access allows the research community to replicate and extend experiments, train and evaluate new models, and publish results without encountering proprietary barriers. As a result, OpenROAD constitutes a natural choice for academic and industrial collaboration, serving as a reproducible testbed in which machine learning and LLM-based agents can be integrated. The availability of its source code and APIs facilitates seamless instrumentation for collecting training data, real-time metrics, and debugging information—all of which are vital for applying advanced ML or agent-based strategies.
@@
In short, leveraging OpenROAD allows us to (1) faithfully represent the full digital implementation flow (floorplanning, placement, clock tree synthesis, routing, etc.), (2) integrate with a community-driven research ecosystem, and (3) avoid opaque “black-box” workflows that impede effective ML experimentation. This environment provides an ideal launchpad for exploring the potential of LLM-based EDA agents that can dynamically interpret intermediate results and fine-tune the flow for improved outcomes.

### v2

@@ IV Experiments and Results
Our experiments use six design instances comprising three circuits (IBEX, AES and JPEG) in two technology nodes (SKY130HD and ASAP7). All experiments are performed on a Google Cloud Platform (GCP) virtual machine with 112 virtual CPUs (C2D AMD Milan) and 220GB RAM. The flow is based on the ce8d36a commit of the OpenROAD-flow-scripts repository [79], along with associated versions of OpenROAD, KLayout, Yosys, and other dependencies.
Although the timeout to run depends on the circuit, each “iteration” is consistently defined as 25 parallel runs, with the exception of the JPEG circuits, which are run with 121212 parallel runs each. Each of these parallel runs uses four vCPUs and 4 GB of RAM. These parallel runs write to different intermediate folders for logs and results, running Yosys and other steps in parallel without reusing intermediate steps. At the end of each iteration, a clean operation moves the logs and folders to a separate directory to avoid contamination.
Our experiments cover two settings. The first, “4-variable” setting involves optimizing over four parameters: Core Utilization, TNS End Percent, LB Add On Place Density, and Clock Period. The first three are in the configuration files of the respective circuits; Clock Period is in the SDC file. The second, “12-variable” setting has eight additional tunable parameters that guide the physical design flow more comprehensively. We distinguish three experimental cases.
@@
  • 12-variable with tool use, which incorporates all 12 variables along with the Bayesian optimization tools.
Our experimental choices reflect the LLM requiring more complex tools to tackle higher-dimensional optimization problems. All experiments involve either 375375375 (4-variable setting) or 600600600 iterations (12-variable setting). The ORFS-agent parameters are configured as follows: (i) the model is claude-3-5-sonnet-20241022; [^1] We do not observe any performance gains from the o1-class of models. (ii) temperature is 0.10.10.1; (iii) nucleus and top-K sampling are disabled; and (iv) to always output 25×4/1225$\times$ 4/1225 × 4 / 12 parameters (the number of parameters the LLM is optimizing over) in the specified format, we utilize JSON Mode (Structured Outputs API). [^2] In terms of API costs, all of our experiments, including all failed runs, cost US \$48. Claude Sonnet 3.5 has since been overtaken by models which are simultaneously better and cheaper by 5 – 10$\times$ per token. As noted above, approaches such as ChatEDA require fine-tuning, which costs significantly more (by at least 10–20$\times$) and requires inference costs for serving the custom model. Appendix B provides additional experimental details.
@@
TABLE I: OR-AutoTuner results. Top block: 375 iterations with 4 parameters. Bottom block: 1000 iterations with 12 parameters. Columns are grouped by optimization: $WL$-only ($WL$, $ECP$*), $ECP$-only ($ECP$, $WL$*), and co-optimization ($WL$**, $ECP$**). Bold font indicates superior results relative to Table II in the corresponding comparison class of (375, 4) or (600, 12) iterations and parameters.
@@
TABLE II: ORFS-agent results for various tuning settings. Top: 375 iterations, 4 parameters (no tool); Middle: 375 iterations, 4 parameters (with tool); Bottom: 600 iterations, 12 parameters (with tool). Each row reports the figures of merit under WL-only, ECP-only, and Co-optimization objectives. Bold font indicates superior results relative to Table I in the corresponding comparison class of (375, 4) or (1000, 12) iterations and parameters.
@@
  • $ECP^{*}$ (resp. $WL^{*}$) denotes the value of $ECP$ (resp. $WL$) when $WL$ (resp. $ECP$) is optimized as the objective; and
  • $ECP^{**}$ and $WL^{**}$ represent the values obtained when both metrics are jointly optimized in a multi-objective setting.
We use OR-AutoTuner as the baseline. We run OR-AutoTuner for 100010001000 iterations with 121212 parameters to optimize, and for 375375375 iterations with 444 parameters to optimize. Results are shown in Table I.
In the following, Section IV-A presents results for single-objective optimization. Then, Subsection IV-B gives results for multi-objective optimization. We then demonstrate ORFS-agent’s capability to perform constrained optimization guided by natural language commands. Appendix C provides additional ablation studies for the ASAP7-IBEX design. Studies on data leakage, model confidence and robustness are presented in Appendix D.
@@ IV-A Single-objective Optimization
There is only a single objective to optimize in this problem. You observe that the value inside the final JSON is $WL$, while the corresponding baseline value is $WL_{\alpha}$. Your effective loss to minimize should be $\frac{WL}{WL_{\alpha}}$.
where $WL_{\alpha}$ denotes the routed wirelength obtained using the default configuration of OpenROAD-flow-scripts and serves as a fixed baseline for normalization. [^5] The baseline results obtained using the default configuration of OpenROAD-flow-scripts are shown in Section IV-C.
However, due to runtime timeouts, typically occurring during the detailed routing stage, the routed wirelength $WL$ may not always be available. In contrast, the estimated wirelength at the end of CTS $WL^{\prime}$ almost always exists. If $WL$ is absent but $WL^{\prime}$ exists, we include the following instruction:
You do not observe the correct value $WL$, but you observe a strong surrogate $WL^{\prime}$, which you should use as a signal. Your effective loss should be $\frac{WL^{\prime}}{WL^{\prime}_{\alpha}}$.
where $WL^{\prime}_{\alpha}$ denotes post-CTS wirelength obtained under the same default configuration. This surrogate-based guidance enables the LLM to continue learning from partial information while maintaining alignment with the optimization objective.
[…]
@@
[…]
  • Comparison between OR-AutoTuner and ORFS-agent: Despite using 40% fewer iterations, ORFS-agent (600 iterations, 12 parameters, with tool use) achieves better results than OR-AutoTuner on 121212 out of 242424 metrics measured for single-objective optimization, thereby exactly drawing on cases of superiority.
@@
[…]
Figure 3: Optimization trajectory of ASAP7-IBEX (co-optimization, 600 iterations, 12 parameters with tool use). Wirelength ($WL$), effective clock period ($ECP$), and instance count (Count) are normalized to baseline values from default ORFS settings.
@@ IV-B Multi-objective Optimization
In the case of multiple losses – e.g., $WL$ and $ECP$ – we instruct the LLM to consider the loss:
$\frac{WL}{WL_{\alpha}}+\frac{ECP}{ECP_{\alpha}}$ (1)
with the same substitution, if applicable, for $ECP^{\prime}$, $WL^{\prime}$ if any of these are unavailable, with the same phrasing adopted above. The experimental results are shown in Table II. Key results are as follows.
[…]
  • No worsening. As with single-objective optimization, ORFS-agent does not worsen any metric – either $ECP$ or wirelength, relative to baselines. In fact, both QoR metrics improve simultaneously.
We visualize the optimization trajectory of instance count alongside wirelength and $ECP$ in Figure 3. We notice that instance count can actually improve for free. To better understand gains on auxiliary metrics (metrics that do not directly enter the optimization function), we examine correlational properties of metrics in Figure 4. We look at how other metrics vary in an optimization run as we carry out ORFS-level runs for ASAP7-IBEX under different parameters and differing optimization goals – namely wirelength, $ECP$, or a combination of the two as defined above. Some metrics do not “come for free” – $ECP$ and power, for instance, have an inverse relationship with each other.
Figure 4: Correlation of metrics for ASAP7-IBEX: wirelength ($WL$), effective clock period ($ECP$), instance area (Area), instance count (Instances) and power (Power). Pearson correlation captures linear relationships; Kendall correlation captures rank-based relationships.
@@ IV-C Baseline Comparisons
We obtain baseline solution metrics using the default ORFS flow parameters seen at the cited commit hash. These metrics are shown in Tables III and  IV. Compared to these baselines, ORFS-agent achieves an improvement of over 13% versus the default ORFS flow parameters, when averaged across all six testcases and single-metric (i.e., $ECP$ and $WL$) comparisons. (For all metrics, lower is better.) For additional context, Table V gives post-route metrics obtained by a proprietary, closed-source tool on ASAP7 circuits; these are quoted from [21].
CTSWL ($WL^{\prime}$) CTSECP ($ECP^{\prime}$) $WL$ $ECP$
@@
TABLE III: Metric values for default ORFS results.
@@
TABLE IV: Derived values for default ORFS results.
@@ IV-D Constrained Optimization under Natural Language Commands
[…]
@@
TABLE VI: Constrained optimization results. Red positive numbers indicate worsening, i.e., an increase relative to baseline; “leeway” indicates the maximum worsening allowed.
@@ V Conclusion and Future Directions
In this work, we propose ORFS-agent, a plug-in open-source framework that integrates LLMs into OpenROAD for hyperparameter optimization. Despite using 40% fewer iterations, ORFS-agent achieves results comparable to OR-AutoTuner. ORFS-agent adopts a modular, provider-agnostic framework, which makes it user-friendly and easy to adopt. ORFS-agent can switch seamlessly among LLM providers (e.g., Anthropic, OpenAI, Gemini) without reconfiguration. Additionally, ORFS-agent avoids the cost and complexity of fine-tuning and deployment, and requires no retraining or infrastructure changes when a new LLM is released. Thus, our ORFS-agent can instantly benefit from model breakthroughs. For example, our preliminary results with Claude 4 Sonnet models (released May 22, 2025) already surpass both Claude 3.5-based ORFS-agent and OR-AutoTuner. Unlike fine-tuning-based approaches, ORFS-agent can rapidly and cost-effectively adopt stronger models, making it both competitive and appealing.
Our future work will explore the following directions. (i) The field has seen the emergence of reasoning-based models, which use a longer chain of thought process before making decisions. These reasoning-based models are then fine-tuned to task-specific agentic models. We aim to integrate such agentic models into ORFS-agent to enhance its decision-making depth and adaptability. (ii) Our current framework uses a fixed set of pre-defined tools. In future work, we will enable the LLM agent to generate its own tools dynamically, which avoids the hand-coding of specific tools and increases flexibility in handling different circuits and technology nodes. (iii) We plan to integrate ORFS-agent with search APIs and RAG tools that enable the agent to access documentation, literature or previous published results. And (iv) we will evaluate ORFS-agent on larger-scale circuits and with more hyperparameters to assess its scalability and robustness. In combination with open-sourcing and OpenROAD integration, we believe that this work will add to the foundations for new research on LLMs for physical design and EDA.
Acknowledgments
This work is partially supported by the Samsung AI Center.
@@ Appendix A EDA Flow and OpenROAD
The EDA flow is highly iterative , and suboptimal outcomes at one step may require going back to earlier steps for re-optimization (e.g., re-synthesizing the gate-level netlist or adjusting the placement). Each step generates logs and metrics that can be integrated into ML pipelines to guide subsequent decisions.
@@
To systematically study LLM- and ML-driven EDA optimizations, we focus on the OpenROAD project [3]. In contrast to commercial EDA offerings, OpenROAD is fully open-source, thus avoiding copyright restrictions and enabling access to the internal details of its place-and-route algorithms, log files, and data structures. This open access allows the researchers to replicate and extend experiments, train and evaluate new models, and publish results without encountering proprietary barriers. [25]
@@
In our work, leveraging OpenROAD allows us to (i) faithfully represent the full digital implementation flow (floorplanning, placement, clock tree synthesis, routing, etc.), (ii) integrate with a community-driven research ecosystem, and (iii) avoid opaque “black-box” workflows that impede effective ML experimentation. In particular, we are able to explore the potential of LLM-based EDA agents that can dynamically interpret intermediate results and fine-tune the flow for improved outcomes.